OpenStudy (anonymous):

Imagine the point (11 -8) on the terminal side of an angle. Find the value of the secant of this angle. i need help asap

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):



OpenStudy (anonymous):

secant is 1/cos

OpenStudy (anonymous):

use a ^2 + b^2 = c^2 to find the third side

OpenStudy (anonymous):

and then you know the hypotenuse to find cos.

OpenStudy (anonymous):

and after finding cos. just do 1/ cos to find secant

OpenStudy (anonymous):

i think this is right but i might be totally wrong

OpenStudy (anonymous):

\[hypotenuse = \sqrt{11^2 + 8^2}\] \[secant=hypotenuse/ initial side\]

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

so i think my answer is sec=sqrt(185)/11?

OpenStudy (anonymous):

yeah!! genius. use calculator to solve it further.

OpenStudy (anonymous):

okay thanks (: sorry it just took me little while to figure it out haha

OpenStudy (anonymous):

np)